# AtomGPT: Atomistic Generative Pre-trained Transformer for Forward and Inverse Materials Design

### v1

@@ Abstract
In Eq. 6, the parameter $\mu^{*}$ is the effective Coulomb potential parameter, which we take as 0.1. It is important to note that the robustness of this workflow was heavily benchmarked against experimental data and higher levels of theory in ref. 30, 53, 54, 55, which indicates that the training data for these deep learning models is of high quality, given the level of theory used to produce the data.
We find a $T_{C}$ of 4.76 K for Nb2B and observe no negative modes in the phonon density of states curve suggesting dynamical stability. Similar DFT calculations can be performed for other structures as we populate more candidates, which can then be added to the JARVIS-DFT Superconductor database 30, 53, 54, 55. Moreover, we can extend this workflow to include other candidate systems, such as nitrides and carbides. As the database of superconductors expands, both forward and inverse models have the potential for further improvement, which is necessary for enhanced predictions. After DFT validation, we can recommend these computationally discovered candidates for experimental synthesis. Superconductor design remains one of the most fascinating areas of research in physics and chemistry. Our work here serves as a proof-of-concept that demonstrates the potential of LLM models, such as GPTs, to accelerate superconductor materials design. Moreover, it can be easily extended to other classes of materials such as alloys, semiconductors, themoelectrics, dielectrics, piezoelectrics, solar cells etc.
@@ Acknowledgements
We thank computational resources from the National Institute of Standards and Technology (NIST). This work was performed with funding from the CHIPS Metrology Program, part of CHIPS for America, National Institute of Standards and Technology, U.S. Department of Commerce. Please note commercial software is identified to specify procedures. Such identification does not imply recommendation by NIST.

### v2

@@ Abstract
In Eq. 6, the parameter $\mu^{*}$ is the effective Coulomb potential parameter, which we take as 0.1. It is important to note that the robustness of this workflow was heavily benchmarked against experimental data and higher levels of theory in ref. 38, 63, 64, 65, which indicates that the training data for these deep learning models is of high quality, given the level of theory used to produce the data.
The $T_{C}$ for Nb2B is predicted to be 4.76 K with no negative modes in the phonon density of states curve suggesting dynamical stability. Similar DFT calculations can be performed for other structures after populating more candidates, which can then be added to the JARVIS-DFT Superconductor database 38, 63, 64, 65. Moreover, we can extend this workflow to include other candidate systems, such as nitrides and carbides. As the database of superconductors expands, both forward and inverse models have the potential for further improvement, which is necessary for enhanced predictions. After DFT validation, we can recommend these computationally discovered candidates for experimental synthesis. Superconductor design remains one of the most fascinating areas of research in physics and chemistry. Our work here serves as a proof-of-concept that demonstrates the potential of LLM models, such as GPTs, to accelerate superconductor materials design. Moreover, it can be easily extended to other classes of materials such as alloys, semiconductors, themoelectrics, dielectrics, piezoelectrics, solar cells and so on. Additionally, AtomGPT’s application of AtomGPT for defect 66 and interface 67 materials design can also be interesting areas of research in the future.
In conclusion, this study introduces AtomGPT, an atomistic generative pre-trained transformer model specifically designed for both forward and inverse materials design. The model is capable of predicting material properties with high accuracy and generating novel material structures with desired properties. By leveraging well-defined structural and chemical features from the JARVIS-Tools, AtomGPT ensures consistent and reliable input generation, enhancing the robustness of property predictions. The separation of forward and inverse models allows for efficient pre-screening and targeted structure generation, respectively. This dual approach facilitates the rapid discovery and optimization of new materials, paving the way for accelerated advancements in materials science. Note that, in this study, I present AtomGPT and large language models (LLMs) not as the ultimate solutions for materials science challenges, but as highly adaptable and tunable tools that can be tailored to a wide range of applications. The versatility of these models allows them to be customized for specific datasets and tasks, making them valuable complements to existing methodologies. By focusing on their adaptability and integration with other tools such as DFT and unified machine learning force-fields, I aim to highlight their potential to enhance and expand the toolkit available to materials scientists, facilitating innovation and discovery across the field. I have demonstrated the application of this workflow by computationally discovering a novel superconductor Nb2B with $T_{C}$ of 4.76 K. These findings highlight the potential of large language models in transforming materials design processes, offering a promising direction for future research and applications in various domains of materials science.
@@ Acknowledgements
K.C. thanks computational resources from the National Institute of Standards and Technology (NIST). This work was performed with funding from the CHIPS Metrology Program, part of CHIPS for America, National Institute of Standards and Technology, U.S. Department of Commerce. Certain commercial equipment, instruments, software, or materials are identified in this paper in order to specify the experimental procedure adequately. Such identifications are not intended to imply recommendation or endorsement by NIST, nor it is intended to imply that the materials or equipment identified are necessarily the best available for the purpose.